OpenStudy (anonymous):

Two trains, the Express and the Commuter, leave the station at the same time. The Express, which heads north, travels 10 km per hour faster than the Commuter, which goes east. If the trains are 100km apart after 2 hours, find the speed of each train.

OpenStudy (anonymous):

((10+k)t)^2 +(kt)^2 plug t=2 ((10+k)2)^2 + 4k^2=10000

OpenStudy (anonymous):

I got k=30

OpenStudy (anonymous):

Solve the following for c\[(2 (10+c))^2+(2c)^2= 100^2\]Commuter's speed = 30 kmh, Express' speed = 40 kmh